OpenStudy (anonymous):

a rectangular pen with a perimeter of 60 ft is needed to house 3 pigs. determine the maximum area of the pen and the dimensions that yield the maximum area. please help! i have this second last unit to cover then i can finally graduate and apply for college i need your help!

OpenStudy (anonymous):

225 ft. 15x15ft

OpenStudy (ranga):

Let L be the length and W be the width of the rectangular pen. Perimeter = 2L + 2W = 60 or L + W = 30 ---- (1) Area A = L * W = L * (30 - L) = 30L - L^2 Complete the square: A = -(L^2 - 30L) = -{ (L - 15)^2 - 225 } If A is plotted against L you will see an inverted parabola with vertex at (15, 225). So the maximum area = 225 ft^2; L = 15 ft. and W = 15 ft.